# Propagation and emission of gravitational waves in the weak-field limit within the Palatini formalism

### v1

@@ Abstract
In the era of gravitational waves physics, when detections of wave fronts are increasing in number, sensibility, frequencies and distances, gravitational physics has entered a period of maximum activity and brilliance. This has open a new window where General Relativity can be challenged in both weak as strong-field regimes. In this paper, we focus on the analysis of gravitational waves propagation and emission in the weak-field regime for gravitational theories within the Palatini formalism. Our results suggest that while propagation in vacuum match General Relativity predictions, emission of gravitational waves from weak sources shows a clear deviation that might be constrained by future detections.
@@ I Introduction
Over the last decades, modifications/extensions of GR have been widely studied, motivated by different phenomena, such analysis has led to a huge increase of knowledge on the way GR might be modified successfully (for a review see Refs. [8, 9, 10]). One of the ways of modifying GR that has been widely explored in the literature is to consider the connection as an independent field [11, 12, 13]. Despite GR is constructed as a geometrical theory that describes pseudo-Riemannian manifolds that are completely characterised by the spacetime metric, non of its principles states that the connection must be the Levi-Civita connection, i.e. that the metricity condition must hold. In addition, GR might be formulated in two different and physically equivalent ways, where the central magnitude is not the curvature but the torsion or the non-metricity [14]. For such theories, known as Teleparallel gravity and Coincident General Relativity respectively, the connection is not given by the Levi-Civita connection, but it is neither let as a free field. However, this has shown that GR might be modified in very different ways. In this paper, we focus on a particular case of the so-called Ricci Based Gravity (RBG) theories, whose starting point is a gravitational action that depends solely on contractions and products of the Ricci tensor, which is constructed by an a priori independent connection [12, 13]. By following the Palatini formalism, which consists on varying the action with respect to the connection as an independent field, the corresponding field equations are obtained for the connection. As far as the gravitational action turns out the Hilbert-Einstein action, the connection becomes the Levi-Civita connection of GR [15]. One of the main advantages of RBG theories is that the field equations for the spacetime metric remain second order while the equations for the connection leads to a direct solution that establishes a connection compatible to a metric tensor. Then, a particular mapping has been established that leads to the same equations as in GR, but now in terms of other metric tensor related to the spacetime metric a conformal transformation [16].The simplest version of RBG theories is the so-called Palatini $f(R)$ gravity where the action is given by a function of the contraction of the Ricci tensor scalar. The main advantage of Palatini $f(R)$ gravity is that the field equations for the spacetime metric remain second order while the equations for the connection leads to a direct solution that establishes a connection compatible to a metric tensor. These theories have been widely analysed in different frameworks, particularly a big effort has been followed for analysing the cosmological implications within these theories, as cosmological late-time acceleration [17, 18, 19, 20, 21] or inflationary models [22, 23, 24, 25, 26, 27]. Moreover, an intensive study has drawn a lot of attention in solutions of compact objects, since one can construct solutions for regular black holes/wormholes [28, 29, 30, 31, 32] or describe well stellar interiors [33, 34].
@@
The present paper intends to focus on the analysis of GWs propagation and emission in the weak-field regime within Palatini $f(R)$ gravity. In general, scalar-tensor theories such as metric $f(R)$ gravities, lead to additional propagating scalar modes that yield to an additional longitudinal polarisation [35, 36, 37, 38, 39, 40, 41, 42]. However, some modifications of GR, such as the Palatini formalism itself [43, 44, 45] or some extensions of Teleparallel gravity [46] do not contain extra propagating modes. The analysis of GWs in modified gravities is essential to test the limits of GR and finding out the way GR might be extended. In this sense, some tests have been suggested in the literature, specifically for metric $f(R)$ gravities, where constraints on the free parameters of the models can be obtained [47, 48]. Moreover, also along the cosmological evolution some modifications of GR might affect some stochastic GWs backgrounds, specially during early cosmology [49, 50, 51]. In order to focus our analysis on the study of GWs in the weak-field limit for Palatini $f(R)$ gravity, we express this theory as a scalar-tensor theory (Brans-Dicke-like model). GWs have been widely explored within Brans-Dicke theories, both in the strong-field regime [52] as in the weak-field limit (see Refs. [54, 55, 56, 57, 58]). Nevertheless, contrary to usual Brans-Dicke-like theories, in Palatini $f(R)$ gravities, the scalar field is not dynamical, what has led the believe that GWs propagation and generation coincide with GR predictions. Here, we show that while propagating GWs in vacuum coincide with GR, emission of GWs has subtle differences with respect to GR, even in the weak field limit. We also point a way such deviations might be constrained.
@@ II Modified Palatini $f(R)$ gravity
Whereas the curvature scalar for the gravitational action is $\mathcal{R}=g^{\mu\nu}\mathcal{R}_{\mu\nu}(\Gamma)$. Note that for the Hilbert-Einstein action, the Palatini formalism directly leads to GR, since the metricity condition is recovered, as shown below. However, for a non-linear function of $\mathcal{R}$ in (1), the metricity condition does not hold and the connection remains independent in principle. Then, by varying the action (1) with respect to the spacetime metric, the field equations are obtained [11]:
@@
$\mathcal{R}_{\mu\nu}(q)=R_{\mu\nu}(g)+\frac{4}{\Omega^{2}}\nabla_{\mu}\Omega\nabla_{\nu}\Omega-\frac{2}{\Omega}\nabla_{\mu}\nabla_{\nu}\Omega-g_{\mu\nu}\frac{g^{\rho\sigma}}{\Omega^{2}}\nabla_{\rho}\Omega\nabla_{\sigma}\Omega-g_{\mu\nu}\frac{\Box\Omega}{\Omega}\ .$ (7)
@@
$R_{\mu\nu}(g)-\frac{1}{2}g_{\mu\nu}R(g)=\frac{\kappa^{2}}{f_{\mathcal{R}}}T_{\mu\nu}-g_{\mu\nu}\frac{\mathcal{R}f_{\mathcal{R}}-f}{2f_{\mathcal{R}}}-\frac{3}{2f_{\mathcal{R}}^{2}}\left[\nabla_{\mu}f_{\mathcal{R}}\nabla_{\nu}f_{\mathcal{R}}-\frac{1}{2}g_{\nu\mu}\nabla_{\lambda}f_{\mathcal{R}}\nabla^{\lambda}f_{\mathcal{R}}\right]+\frac{1}{f_{\mathcal{R}}}\left[\nabla_{\mu}\nabla_{\nu}f_{\mathcal{R}}-g_{\mu\nu}\Box f_{\mathcal{R}}\right]\ .$ (8)
@@
$R_{\mu\nu}-\frac{1}{2}g_{\mu\nu}R=\frac{\kappa^{2}}{\phi}T_{\mu\nu}-\frac{V(\phi)}{2\phi}g_{\mu\nu}-\frac{3}{2\phi^{2}}\left(\partial_{\mu}\phi\partial_{\nu}\phi-\frac{1}{2}g_{\mu\nu}(\partial\phi)^{2}\right)+\frac{1}{\phi}\left(\nabla_{\mu}\nabla_{\nu}\phi-g_{\mu\nu}\Box\phi\right)\ ,$ (10)
@@ III Linearisation of the field equations
$=\overline{G}_{\mu\nu}+\frac{1}{2}\left(\nabla^{\lambda}\nabla_{\mu}h_{\lambda\nu}+\nabla^{\lambda}\nabla_{\nu}h_{\lambda\mu}-\Box h_{\mu\nu}-\nabla_{\mu}\nabla_{\nu}h+\Box h\overline{g}_{\mu\nu}-\nabla_{\alpha}\nabla_{\beta}h^{\alpha\beta}\overline{g}_{\mu\nu}-\overline{R}h_{\mu\nu}\right),$ (13)
@@
$\frac{1}{2}\left(\nabla^{\lambda}\nabla_{\mu}h_{\lambda\nu}+\nabla^{\lambda}\nabla_{\nu}h_{\lambda\mu}-\Box h_{\mu\nu}-\nabla_{\mu}\nabla_{\nu}h+\Box h\overline{g}_{\mu\nu}-\nabla_{\alpha}\nabla_{\beta}h^{\alpha\beta}\overline{g}_{\mu\nu}-\overline{R}h_{\mu\nu}\right)=$
@@
$-\frac{3}{2\overline{\phi}^{2}}\left[\partial_{\mu}\overline{\phi}\partial_{\nu}\delta\phi+\partial_{\mu}\delta\phi\partial_{\nu}\overline{\phi}-\frac{1}{2}h_{\mu\nu}(\partial\overline{\phi})^{2}-\overline{g}_{\mu\nu}\partial^{\alpha}\partial_{\alpha}\delta\phi\right]+\frac{3}{\overline{\phi}^{2}}\frac{\delta\phi}{\overline{\phi}}\left[\partial_{\mu}\overline{\phi}\partial_{\nu}\overline{\phi}-\frac{1}{2}\overline{g}_{\nu\mu}(\partial\overline{\phi})^{2}\right]$
@@ IV Propagation of gravitational waves in vacuum
$\overline{R}=0\ ,\quad T_{\mu\nu}=0\ ,\quad\overline{\phi}=\phi_{0}\ .$ (17)
@@
$\frac{1}{2}\left(\nabla^{\lambda}\nabla_{\mu}h_{\lambda\nu}+\nabla^{\lambda}\nabla_{\nu}h_{\lambda\mu}-\Box h_{\mu\nu}-\nabla_{\mu}\nabla_{\nu}h+\Box h\eta_{\mu\nu}-\nabla_{\alpha}\nabla_{\beta}h^{\alpha\beta}\eta_{\mu\nu}\right)=\nabla_{\mu}\nabla_{\nu}\frac{\delta\phi}{\phi_{0}}-\eta_{\mu\nu}\Box\frac{\delta\phi}{\phi_{0}}\ .$ (18)
This equation is free from gauge choices and covariant derivatives are actually partial derivatives, as correspond for a flat background metric. In order to simplify the equations (18), we can choose the Lorentz Gauge $\partial^{\alpha}h_{\alpha\beta}=0$ to obtain a wave equation. Note also that the r.h.s. of the equations (18) automatically satisfies the Lorentz gauge. By defining the new tensor $\theta_{\mu\nu}=h_{\mu\nu}-\frac{1}{2}h\eta_{\mu\nu}$ the gravitational wave equation yields:
$\Box\theta_{\mu\nu}=-\nabla_{\mu}\nabla_{\nu}\frac{\delta\phi}{\phi_{0}}+\eta_{\mu\nu}\Box\frac{\delta\phi}{\phi_{0}}\ .$ (19)
@@ V Emission of gravitational waves in the weak-field limit
$\Box\theta_{\mu\nu}=-\frac{2\kappa^{2}}{\phi_{0}}T_{\mu\nu}-\nabla_{\mu}\nabla_{\nu}\frac{\delta\phi}{\phi_{0}}+\eta_{\mu\nu}\Box\frac{\delta\phi}{\phi_{0}}\ ,$ (22)
@@
$\Box\theta_{\mu\nu}=-\frac{2\kappa^{2}}{\phi_{0}}T_{\mu\nu}+\frac{1}{\phi_{0}^{2}V^{\prime\prime}(\phi_{0})}\left(\nabla_{\mu}\nabla_{\nu}T-\eta_{\mu\nu}\Box T\right)\ .$ (24)
Hence, the generation of the GWs does not depend only on the presence of perturbed sources of matter but also on the way the energy-momentum tensor varies, as denoted by the presence of second partial derivatives in the equation (24). However, one would expect that for matter sources varying very slowly, the last two terms of the equation (24) might be neglected, recovering the usual case of GR. Nevertheless, in order to keep the analysis as general as possible and study the possible deviations from GR, we are considering that the last two terms might be the same order as the energy-momentum tensor perturbation itself. To solve the equation (24), we define the effective energy momentum tensor as:
$T_{\mu\nu}^{\text{eff}}=\frac{1}{\phi_{0}}T_{\mu\nu}-\frac{1}{2\kappa^{2}\phi_{0}^{2}V^{\prime\prime}(\phi_{0})}\left(\nabla_{\mu}\nabla_{\nu}T-\eta_{\mu\nu}\Box T\right)\ ,$ (25)
@@
Here $\vec{x}$ are spatial vectors, $x^{0}=t$ and $x_{ret}^{0}=t_{ret}=t-|\vec{x}-\vec{x}^{\prime}|$ is the retarded time and we are assuming $c=1$. By considering an enough large distance $|\vec{x}|$ from the source, we can approximate $|\vec{x}-\vec{x}^{\prime}|=r-\vec{x}^{\prime}\cdot\hat{n}$, where $r=|\vec{x}|$ and $\hat{n}$ is the unit vector pointing to the direction of the GW propagation. For simplicity, we work in the transverse-traceless gauge, which by following the standard notation, it is given by:
@@
Then, by inserting the Green function (29) in the expression (27) and integrating over $x^{0\prime}$, it yields:
@@
The aim now is to perform a multipole expansion and show the dependence of the quadrupole moment for this class of theories. To do so, let’s expand the energy-momentum tensor. For enough large distance, we can make the following expansion:
$T_{kl}^{\text{eff}}(t-\frac{r}{c}+\frac{\vec{x}^{\prime}\cdot\hat{n}}{c},x^{\prime})=T_{kl}^{\text{eff}}(t-r,x^{\prime})+\frac{\partial T_{kl}^{\text{eff}}}{dt}\frac{\vec{x}^{\prime}\cdot\hat{n}}{c}+\frac{\partial^{2}T_{kl}^{\text{eff}}}{dt^{2}}\frac{(\vec{x}^{\prime}\cdot\hat{n})(\vec{x}^{\prime}\cdot\hat{n})}{2c^{2}}+\mathcal{O}\left(\frac{1}{c^{3}}\right)\ ,$ (33)
where we have recovered the speed of light to show that the expansion is inversely proportional to $c$. Hence, the expression (32) becomes:
$\theta_{ij}^{TT}=\frac{4G}{r}\Lambda_{ijkl}\left(S^{kl}+\frac{1}{c}n_{m}\dot{S}^{klm}+\frac{1}{2c^{2}}n_{m}n_{p}\ddot{S}^{klmp}+...\right)\ .$ (34)
@@
$S^{kl}$ $=$ $\int d^{3}x^{\prime}T_{\text{eff}}^{km}\delta_{m}^{l}=\int d^{3}x^{\prime}\frac{\partial}{\partial x^{m\prime}}\left(T_{\text{eff}}^{km}x^{l\prime}\right)-\int d^{3}x^{\prime}\frac{\partial T_{\text{eff}}^{km}}{\partial x^{m\prime}}x^{l\prime}=$ (37)
$=$ $-\int d^{3}x^{\prime}\frac{\partial T_{\text{eff}}^{km}}{\partial x^{m\prime}}x^{l\prime}=\int d^{3}x^{\prime}\dot{T}_{\text{eff}}^{k0}x^{l\prime}=\int d^{3}x^{\prime}\ddot{T}_{\text{eff}}^{00}x^{k\prime}x^{l\prime}\ .$ (38)
@@
$S^{kl}=\frac{1}{\phi_{0}}\frac{\partial^{2}}{\partial t^{2}}\int d^{3}x^{\prime}\left(T^{00}-\frac{1}{\phi_{0}V_{0}^{\prime\prime}}\partial_{j}\partial^{j}T\right)x^{k\prime}x^{l\prime}=\ddot{Q}^{kl}\ ,$ (39)

### v2

@@ Abstract
In the era of gravitational waves physics, when detections of wave fronts are increasing in number, sensibility, frequencies and distances, gravitational physics has entered a period of maximum activity and brilliance. This has open a new window where General Relativity can be challenged in both weak as strong-field regimes. In this paper, we focus on the analysis of gravitational waves propagation and emission in the weak-field regime for gravitational theories within the Palatini formalism. Our results show that gravitational waves propagation in vacuum matches General Relativity predictions as well as the multipolar expansion when considering weak sources, independently of the form of the gravitational action.
@@ I Introduction
Over the last decades, modifications/extensions of GR have been widely studied, motivated by different phenomena, such analysis has led to a huge increase of knowledge on the way GR might be modified successfully (for a review see Refs. [8, 9, 10]). One of the ways of modifying GR that has been widely explored in the literature is to consider the connection as an independent field [11, 12, 13]. Despite GR is constructed as a geometrical theory that describes pseudo-Riemannian manifolds that are completely characterised by the spacetime metric, non of its principles states that the connection must be the Levi-Civita connection, i.e. that the metricity condition must hold. In addition, GR might be formulated in two different and physically equivalent ways, where the central magnitude is not the curvature but the torsion or the non-metricity [14]. For such theories, known as Teleparallel gravity and Coincident General Relativity respectively, the connection is not given by the Levi-Civita connection, but it is neither let as a free field. However, this has shown that GR might be modified in very different ways. In this paper, we focus on a particular case of the so-called Ricci Based Gravity (RBG) theories, whose starting point is a gravitational action that depends solely on contractions and products of the Ricci tensor, which is constructed by an a priori independent connection [12, 13]. By following the Palatini formalism, which consists on varying the action with respect to the connection as an independent field, the corresponding field equations are obtained for the connection. As far as the gravitational action turns out the Hilbert-Einstein action, the connection becomes the Levi-Civita connection of GR [15]. One of the main advantages of RBG theories is that the field equations for the spacetime metric remain second order while the equations for the connection leads to a direct solution that establishes a connection compatible to a metric tensor. Then, a particular mapping has been established that leads to the same equations as in GR, but now in terms of other metric tensor related to the spacetime metric a conformal transformation [16].The simplest version of RBG theories is the so-called Palatini $f(R)$ gravity where the action is given by a function of the contraction of the Ricci tensor scalar. The main advantage of Palatini $f(R)$ gravity is that the field equations for the spacetime metric remain second order while the equations for the connection leads to a direct solution that establishes a connection compatible to a metric tensor. These theories have been widely analysed in different frameworks, particularly a big effort has been followed for analysing the cosmological implications within these theories, as cosmological late-time acceleration [17, 18, 19, 20, 21] or inflationary models [22, 23, 24, 25, 26, 27, 28, 29]. Moreover, an intensive study has drawn a lot of attention in solutions of compact objects, since one can construct solutions for regular black holes/wormholes [30, 31, 32, 33, 34] or describe well stellar interiors [35, 36].
@@
The present paper intends to focus on the analysis of GWs propagation and emission in the weak-field regime within Palatini $f(R)$ gravity. In general, scalar-tensor theories such as metric $f(R)$ gravities, lead to additional propagating scalar modes that yield to an additional longitudinal polarisation [37, 38, 39, 40, 41, 42, 43, 44]. However, some modifications of GR, such as the Palatini formalism itself [45, 46, 47] or some extensions of Teleparallel gravity [48] do not contain extra propagating modes. So far, GWs properties in certain theories of modified gravities have been found by using the post-Newtonian formalism [49, 50, 51, 52] and the WKB approach [53]. The analysis of GWs in modified gravities is essential to test the limits of GR and to find out the way GR might be extended. In this sense, some tests have been suggested in the literature, specially for metric $f(R)$ gravities where constraints on the free parameters of the models can be obtained [54, 55]. Moreover, some modifications of GR might affect stochastic GWs backgrounds, specially during early cosmology [56, 57, 58]. In order to focus our analysis on the study of GWs in the weak-field limit for Palatini $f(R)$ gravity, we express this theory as a scalar-tensor theory (Brans-Dicke-like model). GWs have been widely explored within Brans-Dicke theories, both in the strong-field regime [59] as in the weak-field limit (see Refs. [61, 62, 63, 64, 65]). Nevertheless, contrary to usual Brans-Dicke-like theories, the scalar field is not dynamical in Palatini $f(R)$ gravities. Then, by following such procedure we show that propagating GWs in vacuum coincide with GR predictions, as shown previously in [53] by following a different approach. Moreover, by considering weak sources, the corresponding multipole expansion is performed, where the quadrupole and octupole moments are computed, showing that despite the GWs equation is sourced by an effective energy-momentum tensor that includes derivatives of the trace of the energy-momentum tensor, each order of the multipole expansion reduce to the ones predicted by GR.
@@ II Modified Palatini $f(R)$ gravity
Whereas the curvature scalar for the gravitational action is $\mathcal{R}=g^{\mu\nu}\mathcal{R}_{\mu\nu}(\Gamma)$. The independent connection is not necessarily torsionless a priori, however, only its symmetric part plays a role on the equations of motion (see. [11]). Note that for the Hilbert-Einstein action, the Palatini formalism directly leads to GR, since the metricity condition is recovered, as shown below. However, for a non-linear function of $\mathcal{R}$ in (1), the metricity condition does not hold and the connection remains independent in principle. Then, by varying the action (1) with respect to the spacetime metric, the field equations are obtained [11]:
@@
$\mathcal{R}_{\mu\nu}(q)=R_{\mu\nu}(g)+\frac{4}{\Omega^{2}}\nabla_{\mu}\Omega\nabla_{\nu}\Omega-\frac{2}{\Omega}\nabla_{\mu}\nabla_{\nu}\Omega$
$-g_{\mu\nu}\frac{g^{\rho\sigma}}{\Omega^{2}}\nabla_{\rho}\Omega\nabla_{\sigma}\Omega-g_{\mu\nu}\frac{\Box\Omega}{\Omega}\ .$ (7)
@@
$R_{\mu\nu}(g)-\frac{1}{2}g_{\mu\nu}R(g)=\frac{\kappa^{2}}{f_{\mathcal{R}}}T_{\mu\nu}-g_{\mu\nu}\frac{\mathcal{R}f_{\mathcal{R}}-f}{2f_{\mathcal{R}}}$
$-\frac{3}{2f_{\mathcal{R}}^{2}}\left[\nabla_{\mu}f_{\mathcal{R}}\nabla_{\nu}f_{\mathcal{R}}-\frac{1}{2}g_{\nu\mu}\nabla_{\lambda}f_{\mathcal{R}}\nabla^{\lambda}f_{\mathcal{R}}\right]$
$+\frac{1}{f_{\mathcal{R}}}\left[\nabla_{\mu}\nabla_{\nu}f_{\mathcal{R}}-g_{\mu\nu}\Box f_{\mathcal{R}}\right]\ .$ (8)
@@
$R_{\mu\nu}-\frac{1}{2}g_{\mu\nu}R=\frac{\kappa^{2}}{\phi}T_{\mu\nu}-\frac{V(\phi)}{2\phi}g_{\mu\nu}$
$-\frac{3}{2\phi^{2}}\left(\partial_{\mu}\phi\partial_{\nu}\phi-\frac{1}{2}g_{\mu\nu}(\partial\phi)^{2}\right)$
$+\frac{1}{\phi}\left(\nabla_{\mu}\nabla_{\nu}\phi-g_{\mu\nu}\Box\phi\right)\ ,$ (10)
@@ III Linearisation of the field equations
$=\overline{G}_{\mu\nu}+\frac{1}{2}\big{(}\nabla^{\lambda}\nabla_{\mu}h_{\lambda\nu}+\nabla^{\lambda}\nabla_{\nu}h_{\lambda\mu}-\Box h_{\mu\nu}$
$-\nabla_{\nu}\nabla_{\mu}h+\Box h\overline{g}_{\mu\nu}-\nabla_{\alpha}\nabla_{\beta}h^{\alpha\beta}\overline{g}_{\mu\nu}-\overline{R}h_{\mu\nu}$
$+\overline{g}_{\mu\nu}h^{\alpha\beta}\overline{R}_{\alpha\beta}\big{)},$ (13)
@@
$\frac{1}{2}\left(\nabla^{\lambda}\nabla_{\mu}h_{\lambda\nu}+\nabla^{\lambda}\nabla_{\nu}h_{\lambda\mu}-\Box h_{\mu\nu}-\nabla_{\nu}\nabla_{\mu}h+\Box h\overline{g}_{\mu\nu}-\nabla_{\alpha}\nabla_{\beta}h^{\alpha\beta}\overline{g}_{\mu\nu}-\overline{R}h_{\mu\nu}+\overline{g}_{\mu\nu}h^{\alpha\beta}\overline{R}_{\alpha\beta}\right)=$
@@
$-\frac{3}{2\overline{\phi}^{2}}\left[\partial_{\mu}\overline{\phi}\partial_{\nu}\delta\phi+\partial_{\mu}\delta\phi\partial_{\nu}\overline{\phi}-\frac{1}{2}h_{\mu\nu}(\partial\overline{\phi})^{2}-\overline{g}_{\mu\nu}\partial^{\alpha}\overline{\phi}\partial_{\alpha}\delta\phi\right]+\frac{3}{\overline{\phi}^{2}}\frac{\delta\phi}{\overline{\phi}}\left[\partial_{\mu}\overline{\phi}\partial_{\nu}\overline{\phi}-\frac{1}{2}\overline{g}_{\nu\mu}(\partial\overline{\phi})^{2}\right]$
@@
$+\overline{g}_{\mu\nu}\frac{\overline{g}^{\lambda\alpha}}{2}(\nabla_{\beta}h_{\lambda\alpha}+\nabla_{\alpha}h_{\lambda\beta}-\nabla_{\lambda}h_{\alpha\beta})\partial_{\beta}\overline{\phi}-h_{\mu\nu}\Box\overline{\phi}-\nabla_{\mu}\nabla_{\nu}\overline{\phi}\frac{\delta\phi}{\overline{\phi}}-\overline{g}_{\mu\nu}\Box\overline{\phi}\frac{\delta\phi}{\overline{\phi}}\Bigg{]}.$ (15)
@@ IV Propagation of gravitational waves in vacuum
$\overline{R}=0\ ,\quad\overline{R}_{\alpha\beta}=0,\quad T_{\mu\nu}=0\ ,\quad\overline{\phi}=\phi_{0}\ .$ (17)
@@
$\frac{1}{2}\big{(}\nabla^{\lambda}\nabla_{\mu}h_{\lambda\nu}+\nabla^{\lambda}\nabla_{\nu}h_{\lambda\mu}-\Box h_{\mu\nu}$
$-\nabla_{\nu}\nabla_{\mu}h+\Box h\eta_{\mu\nu}-\nabla_{\alpha}\nabla_{\beta}h^{\alpha\beta}\eta_{\mu\nu}\big{)}$
$=\nabla_{\mu}\nabla_{\nu}\frac{\delta\phi}{\phi_{0}}-\eta_{\mu\nu}\Box\frac{\delta\phi}{\phi_{0}}\ .$ (18)
This equation is free from gauge choices and covariant derivatives are actually partial derivatives, as correspond for a flat background metric. In order to simplify the equations (18), we can choose the Lorentz Gauge $\partial^{\alpha}\theta_{\alpha\beta}=0$ to obtain the usual wave equation. Note also that the r.h.s. of the equations (18) automatically satisfies the Lorentz gauge. By defining the new tensor $\theta_{\mu\nu}=h_{\mu\nu}-\frac{1}{2}h\eta_{\mu\nu}$ the gravitational wave equation yields:
$\Box\theta_{\mu\nu}=-2\nabla_{\mu}\nabla_{\nu}\frac{\delta\phi}{\phi_{0}}+2\eta_{\mu\nu}\Box\frac{\delta\phi}{\phi_{0}}\ .$ (19)
@@ V Emission of gravitational waves in the weak-field limit
$\Box\theta_{\mu\nu}=-\frac{2\kappa^{2}}{\phi_{0}}T_{\mu\nu}-2\nabla_{\mu}\nabla_{\nu}\frac{\delta\phi}{\phi_{0}}+2\eta_{\mu\nu}\Box\frac{\delta\phi}{\phi_{0}}\ ,$ (22)
@@
$\Box\theta_{\mu\nu}=-\frac{2\kappa^{2}}{\phi_{0}}T_{\mu\nu}$
$+\frac{2}{\phi_{0}^{2}V^{\prime\prime}(\phi_{0})}\left(\nabla_{\mu}\nabla_{\nu}T-\eta_{\mu\nu}\Box T\right)\ .$ (24)
Hence, the generation of the GWs does not depend only on the presence of perturbed sources of matter but also on the way the energy-momentum tensor varies, as denoted by the presence of second partial derivatives in the equation (24). However, one would expect that for matter sources varying very slowly, the last two terms of the equation (24) might be neglected, recovering the usual case of GR. Nevertheless, in order to keep the analysis as general as possible and study the possible deviations from GR, we are considering that the last two terms might be the same order as the energy-momentum tensor perturbation itself. To solve the equation (24), we define the effective energy-momentum tensor as:
$T_{\mu\nu}^{\text{eff}}=\frac{1}{\phi_{0}}T_{\mu\nu}-\frac{1}{\kappa^{2}\phi_{0}^{2}V^{\prime\prime}(\phi_{0})}\left(\nabla_{\mu}\nabla_{\nu}T-\eta_{\mu\nu}\Box T\right)\ ,$ (25)
@@
Here $\vec{x}$ are spatial vectors, $x^{0}=t$ and $x_{ret}^{0}=t_{ret}=t-|\vec{x}-\vec{x}^{\prime}|$ is the retarded time and we are assuming $c=1$. Then, the solution (27) yields:
$\theta_{\mu\nu}=4G\int d^{3}x^{\prime}\frac{1}{|x-x^{\prime}|}T_{\mu\nu}^{\text{eff}}(x^{\prime})\ .$ (30)
By considering an enough large distance $|\vec{x}|$ from the source, we can approximate $|\vec{x}-\vec{x}^{\prime}|=r-\vec{x}^{\prime}\cdot\hat{n}$, where $r=|\vec{x}|$ and $\hat{n}$ is the unit vector pointing to the direction of the GW propagation. Since the volume integral in (30) is taken above the source, the transverse-traceless (TT) gauge can be applied, which by following the standard notation, it is given by:
@@
Then, the integral (30) leads to:
@@
Note that inside the source, the TT gauge can not be applied and the following calculations will not be valid what implies important differences, as pointed at the end of the paper. Nevertheless, the aim now is to perform a multipole expansion and show the dependence of the quadrupole moment for this class of theories. To do so, let’s expand the energy-momentum tensor. For enough large distance, we can make the following expansion:
$T_{kl}^{\text{eff}}(t-\frac{r}{c}+\frac{\vec{x}^{\prime}\cdot\hat{n}}{c},x^{\prime})=T_{kl}^{\text{eff}}(t-r,x^{\prime})$
$+\frac{\partial T_{kl}^{\text{eff}}}{\partial t}\frac{(\vec{x}^{\prime}\cdot\hat{n})}{c}+\frac{\partial^{2}T_{kl}^{\text{eff}}}{\partial t^{2}}\frac{(\vec{x}^{\prime}\cdot\hat{n})(\vec{x}^{\prime}\cdot\hat{n})}{2c^{2}}+\mathcal{O}\left(\frac{1}{c^{3}}\right)\ ,$ (34)
where we have recovered the speed of light to show that the expansion is inversely proportional to powers of $c$. Hence, the expression (33) becomes:
$\theta_{ij}^{TT}=\frac{4G}{rc^{4}}\Lambda_{ijkl}\bigg{(}S^{kl}+\frac{1}{c}n_{m}\dot{S}^{klm}+$
$\frac{1}{2c^{2}}n_{m}n_{p}\ddot{S}^{klmp}+...\bigg{)}\ .$ (35)
@@
$S^{kl}$ $=$ $\int d^{3}x^{\prime}T_{\text{eff}}^{km}\delta_{m}^{l}$ (38)
$=$ $\int d^{3}x^{\prime}\frac{\partial}{\partial x^{m\prime}}\left(T_{\text{eff}}^{km}x^{l\prime}\right)-\int d^{3}x^{\prime}\frac{\partial T_{\text{eff}}^{km}}{\partial x^{m\prime}}x^{l\prime}=$
$=$ $-\int d^{3}x^{\prime}\frac{\partial T_{\text{eff}}^{km}}{\partial x^{m\prime}}x^{l\prime}=\int d^{3}x^{\prime}\dot{T}_{\text{eff}}^{k0}x^{l\prime}$
$=$ $\int d^{3}x^{\prime}\ddot{T}_{\text{eff}}^{00}x^{k\prime}x^{l\prime}\ .$
@@
$S^{kl}$ $=\frac{1}{\phi_{0}}\frac{\partial^{2}}{\partial t^{2}}\int d^{3}x^{\prime}\left(T^{00}-\frac{1}{\kappa^{2}\phi_{0}V_{0}^{\prime\prime}}\partial_{j}\partial^{j}T\right)x^{k\prime}x^{l\prime}$
$=\ddot{Q}^{kl}\ ,$ (39)